OpenStudy (faithspring):

solve by factoring 3. x2 – 16 = 0

alones (alones):

Is 3 part of a problem?

OpenStudy (fortytherapper):

I think it's just x^2 - 16 =0

jhonyy9 (jhonyy9):

x^2 -16 =0 this is a different of square x^2 -16 = (x-4)(x+4) hope this will help you